### v2

@@ Abstract
Results: When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 85% reduction or a human crash rate that is 6.7 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Police-reported and any-injury-reported crashed vehicle rate reductions for the ADS were statistically significant when compared in San Francisco and Phoenix, as well as combined across all locations. The any property damage or injury comparison had statistically significant decrease in 3 comparisons, but also non-significant results in 3 other benchmarks. When excluding ADS crashes with a delta-V less than 1 mph (a measure of sensitivity to lower reporting threshold), about half of the ADS collisions were excluded resulting in comparisons that showed a large statistically significant reduction in all comparisons except for one comparison from San Francisco.
@@ RESULTS
Table 7 shows the ADS to Human rate ratio and 95% confidence intervals for police-reported and any-injury-reported crashed vehicle rates. The ADS had a lower police-reported and Blincoe-adjusted any-injury-reported crashed vehicle rate that was statistically significant in Phoenix, San Francisco, and when considered in aggregate. The observed any-injury-reported comparison was not statistically significant in Phoenix and when compared to a national average, but significant in San Francisco and in the mileage blended comparison. The comparison of the ADS to the observed any-injury-reported benchmark serves as an analysis of the sensitivity of the results to the Blincoe injury underreporting estimate. The results show that even when not accounting for known underreporting in human data, the point estimates still show the ADS had a lower any-injury-reported crash rate, although not all comparisons were statistically significant due to the low event counts and mileage. Inverting the rate ratios in Table 7, humans have between 1.9 and 3.4 times the rate of police-reported crashes and between 4.2 and 10 times the rate of Blincoe-adjusted any-injury-reported crashes as the Waymo ADS. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 85% reduction or a human crash rate that is 6.7 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Figure 3 shows the percent reduction of the ADS for police-reported and any-injury-reported crashed vehicle rates and 95% confidence intervals. This percent reduction is another common form of the rate ratios shown in Table 7.
@@
Any-Injury-Reported Scanlon et al. (2023) Observed Any-Injury-Reported Phoenix 1.24 0.4 0.30 0.03 1.23
@@
Total Mileage Blended 1.91 0.4 0.22* 0.03 0.71
Total National Average 1.21 0.4 0.35 0.05 1.13
Scanlon et al. (2023) Blincoe-adjusted Any-Injury-Reported Phoenix 1.81 0.4 0.21* 0.02 0.84
@@
Total Mileage Blended 2.80 0.4 0.15* 0.02 0.49
Total National Average 1.76 0.4 0.24* 0.04 0.77
@@ DISCUSSION
The benchmarks and ADS data presented in this paper are overall crash involvement rates which do not consider crash contribution of those involved. Di Lillo et al. (2023) compared third party liability property damage and bodily injury insurance claims rates between the Waymo ADS operations and a human benchmark. A liability claim is a request for compensation when someone is responsible for damage to property or injury to another person, typically following a collision. In Di Lillo et al. (2023), only 3rd party liability claims that were resolved or were likely to be resolved with a payment were included as these types of claims are a proxy for responsibility. Di Lillo et al. (2023) found that over 3.8 million miles of RO driving, the Waymo RO service had a statistically significant, lower property damage and bodily injury claims rate compared to a human benchmark based on aggregate human insurance claims data restricted to zip codes where the Waymo service operates. Examining third party liability claims allows an analysis of the degree to which the ADS contributes to crash causation, that is, the ADS safety impact on other road users. The results of the current study, in contrast, examine the overall crashed vehicle rate regardless of the contribution to the crash’s causation. Many of the crashes in the RO dataset examined in this paper featured the Waymo ADS parked or stopped appropriately at a traffic control device where the ADS’s behavior had little or no contribution to the crash’s cause. Examining ADS performance from both a crash causation and overall crash rate is important because it provides assessment of (a) performance in crashes where there was some contribution (3rd party liability claims), (b) performance in crashes where there was no contribution, and (c) the combination (overall crash rates). Di Lillo et al. (2023) reported that Waymo RO operations reduced property damage liability claims by 76% and bodily injury liability claims by 100%. This magnitude of claims reduction is similar to the magnitude in police-reported (49% to 71% reduction) and any-injury-reported (76% and 90% reduction) reductions reported in the current study, which did not consider contribution to causing the crash.
@@ CONCLUSIONS
This study compared the crashed vehicle rate of the Waymo RO service to human benchmarks for any property damage or injury, police-reported, and any-injury-reported outcomes. The results show that the Waymo RO service had a lower police-reported and any-injury-reported crashed vehicle rate compared to the benchmarks. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 85% reduction or a human crash rate that is 6.7 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Because the police-reported and any-injury-reported crashed ADS vehicle rates were much lower than the benchmark rates, the ADS any-injury-reported and police-reported crashed vehicle rates were statistically significant over a total of 7.14 million RO miles when compared individually in Phoenix and San Francisco and when compared overall. Although these results are promising, they are based on a small number of ADS events (15 police-reported and 3 any injury-reported). Thus, these comparisons may be subject to statistical noise as more miles and events are observed in the future. The comparison for Los Angeles was not statistically significant for zero observed ADS events but few miles. The results for the police-reported and any-injury-reported outcome groups, however, show a positive safety impact of the ADS in these types of crashes.
@@ A.1 Selection of SGO Events
Police-reported status was determined using a combination of the NHTSA SGO field “Law Enforcement Investigating?” and internal Waymo records. There were 10 SGO-reported crashes that indicated “Yes” for the “Law Enforcement Investigating?” field. The two (2) SGO-reported crashes with an “Unknown” reported in the “Law Enforcement Investigating?” field also likely had a police report generated based on review of Waymo operations records. Additionally, in two events (30270-6548 and 30270-1583), the SGO field “Law Enforcement Investigating?” is “No”, but operational records indicate a police report may have been filed. These two events were also conservatively included in the SGO police-reported category. Event #1 from Victor et al. (2023) occurred before the NHTSA SGO-reporting period but was reported to police based on review of operational records. To determine any-injury-reported crashes, the “Highest Injury Severity Alleged” SGO field was used. Any crash that did not have “No Injuries Reported” was considered an any-injury-reported crash. In this dataset, there were 3 crashes with “Minor” injuries reported. Crashes where the ADS was not-in-transport were also excluded from police-reported and alleged any-injury-reported crashes. All ADS not-in-transport crashes had no reported law enforcement investigation and no injuries reported as part of the SGO data.
@@
30270-6149 NA PHX True True True True False
@@
As stated in the Methods section (section METHODS), the “NHTSA SGO In-Transport Exclude Low delta-V (DV)” category was derived by applying a impulse-momentum collision model for vehicle-to-vehicle crashes. The collision model used was an impulse-momentum model using the inertial properties of the ADS equipped vehicle and estimated inertial properties of the crash partner, as described in Scanlon et al. (2021). Each of the 7 crashes with fixed or non-fixed objects was examined individually to estimate a delta-V. Of the 7 crashes with fixed or non-fixed objects, 5 were excluded for having a low delta-V. Crashes where the ADS was not-in-transport were also excluded from this group. Three (3) fixed object crashes resulted in underbody damage when the Waymo vehicle “bottomed out” entering a parking lot driveway (30270-5612, 30270-5611) or tire damage after driving through a pothole (30270-5114) which resulted in less than the 1.0 mph delta-V threshold. Two (2) crashes involved non-fixed objects with low mass (<100 kg): a cardboard box (30270-5081) and a swinging parking lot gate that was not fixed (30270-4363). The single crash involving a cyclist (30270-5456) was included in this group.
@@ 14 Rate Ratio and (lower, upper) confidence intervals
15 ("Any-injury PHX", 1.81, phx_benchmark_m_miles, 2, phx_waymo_m_miles),
@@
6Comparison: Any-injury PHX, Rate Ratio: 0.21, CI lower: 0.02, CI upper: 0.84

### v3

@@ Abstract
Results: When considering all locations together, the any-injury-reported crashed vehicle rate was 0.6 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 80% reduction or a human crash rate that is 5 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate.
Police-reported crashed vehicle rate reductions for the ADS were statistically significant when compared in San Francisco and Phoenix, as well as combined across all locations and the any-injury-reported reductions were statistically significant in San Francisco and in all locations. The any property damage or injury comparison had statistically significant decrease in 3 comparisons, but also non-significant results in 3 other benchmarks. When excluding ADS crashes with a delta-V less than 1 mph (a measure of sensitivity to lower reporting threshold), about half of the ADS collisions were excluded resulting in comparisons that showed a large statistically significant reduction in all comparisons except for one comparison from San Francisco.
@@ RESULTS
Table 7 shows the ADS to Human rate ratio and 95% confidence intervals for police-reported and any-injury-reported crashed vehicle rates. The ADS had a lower police-reported and Blincoe-adjusted any-injury-reported crashed vehicle rate. The comparison was statistically significant in all locations for the police-reported outcome and in San Francisco and when considered in aggregate for the any-injury-reported outcome. The observed any-injury-reported comparison was not statistically significant in Phoenix and when compared to a national average, but significant in San Francisco and in the mileage blended comparison. The comparison of the ADS to the observed any-injury-reported benchmark serves as an analysis of the sensitivity of the results to the Blincoe injury underreporting estimate. The results show that even when not accounting for known underreporting in human data, the point estimates still show the ADS had a lower any-injury-reported crash rate, although not all comparisons were statistically significant due to the low event counts and mileage. Inverting the rate ratios in Table 7, humans have between 1.9 and 3.4 times the rate of police-reported crashes and between 3.1 and 10 times the rate of Blincoe-adjusted any-injury-reported crashes as the Waymo ADS. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.6 incidents per million miles (IPMM) for the ADS vs 2.8 IPMM for the human benchmark, an 80% reduction or a human crash rate that is 5 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Figure 3 shows the percent reduction of the ADS for police-reported and any-injury-reported crashed vehicle rates and 95% confidence intervals. This percent reduction is another common form of the rate ratios shown in Table 7.
@@
Any-Injury-Reported Scanlon et al. (2023) Observed Any-Injury-Reported Phoenix 1.24 0.6 0.45 0.07 1.47
@@
Total Mileage Blended 1.91 0.6 0.29* 0.06 0.82
Total National Average 1.21 0.6 0.46 0.10 1.30
Scanlon et al. (2023) Blincoe-adjusted Any-Injury-Reported Phoenix 1.81 0.6 0.31 0.05 1.01
@@
Total Mileage Blended 2.80 0.6 0.20* 0.04 0.56
Total National Average 1.76 0.6 0.32* 0.07 0.90
@@ DISCUSSION
The benchmarks and ADS data presented in this paper are overall crash involvement rates which do not consider crash contribution of those involved. Di Lillo et al. (2023) compared third party liability property damage and bodily injury insurance claims rates between the Waymo ADS operations and a human benchmark. A liability claim is a request for compensation when someone is responsible for damage to property or injury to another person, typically following a collision. In Di Lillo et al. (2023), only 3rd party liability claims that were resolved or were likely to be resolved with a payment were included as these types of claims are a proxy for responsibility. Di Lillo et al. (2023) found that over 3.8 million miles of RO driving, the Waymo RO service had a statistically significant, lower property damage and bodily injury claims rate compared to a human benchmark based on aggregate human insurance claims data restricted to zip codes where the Waymo service operates. Examining third party liability claims allows an analysis of the degree to which the ADS contributes to crash causation, that is, the ADS safety impact on other road users. The results of the current study, in contrast, examine the overall crashed vehicle rate regardless of the contribution to the crash’s causation. Many of the crashes in the RO dataset examined in this paper featured the Waymo ADS parked or stopped appropriately at a traffic control device where the ADS’s behavior had little or no contribution to the crash’s cause. Examining ADS performance from both a crash causation and overall crash rate is important because it provides assessment of (a) performance in crashes where there was some contribution (3rd party liability claims), (b) performance in crashes where there was no contribution, and (c) the combination (overall crash rates). Di Lillo et al. (2023) reported that Waymo RO operations reduced property damage liability claims by 76% and bodily injury liability claims by 100%. This magnitude of claims reduction is similar to the magnitude in police-reported (49% to 71% reduction) and any-injury-reported (69% and 90% reduction) reductions reported in the current study, which did not consider contribution to causing the crash.
@@ CONCLUSIONS
This study compared the crashed vehicle rate of the Waymo RO service to human benchmarks for any property damage or injury, police-reported, and any-injury-reported outcomes. The results show that the Waymo RO service had a lower police-reported and any-injury-reported crashed vehicle rate compared to the benchmarks. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.6 incidents per million miles (IPMM) for the ADS vs 2.8 IPMM for the human benchmark, an 80% reduction or a human crash rate that is 5 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. The ADS any-injury-reported and police-reported crashed vehicle rates lower than the benchmark rates at a statistically significant level over a total of 7.14 million RO miles when compared individually in Phoenix and San Francisco and when compared overall (except for the any-injury-reported comparison in Phoenix). Although these results are promising, they are based on a small number of ADS events (15 police-reported and 4 any injury-reported). Thus, these comparisons may be subject to statistical noise as more miles and events are observed in the future. The comparison for Los Angeles was not statistically significant for zero observed ADS events but few miles. The results for the police-reported and any-injury-reported outcome groups, however, show a positive safety impact of the ADS in these types of crashes.
@@ A.1 Selection of SGO Events
Police-reported status was determined using a combination of the NHTSA SGO field “Law Enforcement Investigating?” and internal Waymo records. There were 10 SGO-reported crashes that indicated “Yes” for the “Law Enforcement Investigating?” field. The two (2) SGO-reported crashes with an “Unknown” reported in the “Law Enforcement Investigating?” field also likely had a police report generated based on review of Waymo operations records. Additionally, in two events (30270-6548 and 30270-1583), the SGO field “Law Enforcement Investigating?” is “No”, but operational records indicate a police report may have been filed. These two events were also conservatively included in the SGO police-reported category. Event #1 from Victor et al. (2023) occurred before the NHTSA SGO-reporting period but was reported to police based on review of operational records. To determine any-injury-reported crashes, the “Highest Injury Severity Alleged” SGO field was used. Any crash with “Minor”, “Moderate”, “Serious”, or “Fatality” as the maximum severity was considered an any-injury-reported crash. In this dataset, there were 3 crashes with “Minor” injuries reported. In addition, there was one crash (30270-6149) with “Unknown” maximum severity. The written SGO case narrative for this case states there were reported injuries and a person involved in the crash “was transported from the scene to a hospital for medical treatment.” This crash with “Unknown” severity was considered an any-injury-reported crash in this study, bringing to total injury crash count to 4. Other crashes with “Unknown” maximum severity due to missing information (e.g., due to another party leaving the scene without exchanging information) were excluded from the any-injury-reported outcome, as these types of missing data are also excluded from the human crash databases used to derived the benchmarks used in this study (Scanlon et al., 2023). Crashes where the ADS was not-in-transport were also excluded from police-reported and alleged any-injury-reported crashes. All ADS not-in-transport crashes had no reported law enforcement investigation and no injuries reported as part of the SGO data.
@@
30270-6149 NA PHX True True True True True
@@
As stated in the Methods section, the “NHTSA SGO In-Transport Exclude Low delta-V (DV)” category was derived by applying a impulse-momentum collision model for vehicle-to-vehicle crashes. The collision model used was an impulse-momentum model using the inertial properties of the ADS equipped vehicle and estimated inertial properties of the crash partner, as described in Scanlon et al. (2021). Each of the 7 crashes with fixed or non-fixed objects was examined individually to estimate a delta-V. Of the 7 crashes with fixed or non-fixed objects, 5 were excluded for having a low delta-V. Crashes where the ADS was not-in-transport were also excluded from this group. Three (3) fixed object crashes resulted in underbody damage when the Waymo vehicle “bottomed out” entering a parking lot driveway (30270-5612, 30270-5611) or tire damage after driving through a pothole (30270-5114) which resulted in less than the 1.0 mph delta-V threshold. Two (2) crashes involved non-fixed objects with low mass (<100 kg): a cardboard box (30270-5081) and a swinging parking lot gate that was not fixed (30270-4363). The single crash involving a cyclist (30270-5456) was included in this group.
@@ 14 Rate Ratio and (lower, upper) confidence intervals
15 ("Any-injury PHX", 1.81, phx_benchmark_m_miles, 3, phx_waymo_m_miles),
@@
6Comparison: Any-injury PHX, Rate Ratio: 0.31, CI lower: 0.05, CI upper: 1.01